OpenStudy (anonymous):

I am stuck on factoring this X^3-x^2-x+1 I got X^2(x-1)

OpenStudy (agent47):

x^2(x-1)-x+1=x^2(x-1)-(x-1)=(x-1)[x^2-1]=(x+1)(x-1)^2

OpenStudy (anonymous):

@precal

OpenStudy (anonymous):

You lost me

OpenStudy (precal):



OpenStudy (precal):

agent 47 was correct but don't forget x^2 -1 is a difference of squares

OpenStudy (precal):



OpenStudy (precal):

hope that helps